# UAV-DETR: Efficient End-to-End Object Detection for Unmanned Aerial Vehicle Imagery

### v1

@@ IV-C Ablation Studies
The baseline RT-DETR-R18 achieves an AP of 26.7 and AP50 of 44.6. After incorporating Inner-SIoU, AP increased to 27.1, showing that improving the loss function positively impacts performance. Adding the MSFF-FE module further improves the AP to 28.4, demonstrating the benefits of incorporating multi-scale feature fusion and frequency enhancement. The addition of FD improves AP50, reaching 47.1. Incorporating the SAC module results in further gains, with AP reaching 28.9 and AP50 improving to 47.7. When all components are combined, UAV-DETR-R18 achieves the highest performance with an AP of 29.8 and an AP50 of 48.8, showcasing the cumulative impact of each module on detection accuracy. As shown in Table IV, our experiments demonstrate that setting the ratio of Inner-SIOU to 1.25 is an appropriate choice. Lastly, we calculated the Frames Per Second (FPS) for both the baseline and UAV-DETR models using the PyTorch implementation in 32-bit Floating Point Precision, as shown in Table V. The results show that UAV-DETR is capable of meeting real-time requirements.
@@ IV-D Visualization
In Fig. 5, we present heatmaps for small objects in the VisDrone dataset, focusing on backpropagation through the bounding box predictions. Compared to the baseline model, UAV-DETR demonstrates a notably improved ability to localize small objects. In the heatmaps of our model, small objects exhibit higher heat values, indicating that the model can more effectively capture the features of these small objects. Additionally, it can be observed that UAV-DETR pays more attention to the surrounding information of small objects, demonstrating the model’s ability to better utilize contextual information during detection. As a result, as shown in the yellow box in Fig. 5, UAV-DETR also performs well in localizing occluded objects.
@@ IV-E Discussion
There are several reasons that contribute to this improved performance. Firstly, our model retains more high-frequency features, which are crucial for detecting small objects. In traditional feature fusion and down-sampling processes, high-frequency features are often lost, making it harder to recover edge and texture details. These details are particularly important for UAV-OD. To address this, we introduced the MSFF-FE and FD modules, which enable the model to combine spatial and frequency domain information during both feature fusion and down-sampling, ensuring that important high-frequency components are preserved.
Secondly, our model makes better use of contextual information. When small objects are difficult to detect based on fine details, their surrounding semantic context becomes crucial. The operations in the frequency domain help the model capture global patterns, improving detection accuracy. However, these operations can sometimes lead to misalignment between the semantic and spatial information of different feature maps. To address this, we designed the SAC module, which aligns features from different fusion paths and enhances overall detection performance. The ablation study in Table III supports the effectiveness of these modules.
@@ V Conclusion
We design UAV-DETR, a real-time end-to-end object detector specifically designed for UAV imagery. By introducing the MSFF-FE module, FD module, and SAC module, UAV-DETR helps alleviate the difficulties of detecting small and occluded objects in aerial images. Experimental results on the VisDrone and UAVVaste datasets demonstrate that our method achieves higher accuracy than existing approaches with similar computational costs, all while maintaining real-time inference speeds. Future work will focus on improving its robustness to noise.

### v3

@@ IV-D Visualization
In Fig. 5, we present heatmaps for small objects in the VisDrone dataset, focusing on backpropagation through the predictions of the bounding box. Compared to the baseline model, UAV-DETR demonstrates a notably improved ability to localize small objects. In the heatmaps of our model, small objects exhibit higher heat values, which indicates that the model can more effectively capture the features of these small objects. Additionally, it can be observed that UAV-DETR pays more attention to the surrounding information of small objects, which demonstrates the model’s ability to better utilize contextual information during detection. As a result, as shown in the yellow box in Fig. 5, UAV-DETR also performs well in localizing occluded objects. Therefore, UAV-DETR does not severely degrade the baseline’s real-time performance. After deployment on different hardware platforms, FPS can obtain significant improvements[36].
@@ IV-E Discussion
There are several reasons that contribute to this improved performance. Firstly, our model retains more high-frequency features, which are crucial for detecting small objects. In traditional feature fusion and downsampling processes, high-frequency features are often lost, which makes it harder to recover edge and texture details. These details are particularly important for UAV-OD. To address this, we introduced the MSFF-FE and FD modules, which enable the model to combine spatial and frequency domain information during feature fusion and downsampling to eusure that important high-frequency components are preserved.
Secondly, our model makes better use of contextual information. When small objects are difficult to detect based on fine details, their surrounding semantic context becomes crucial. The operations in the frequency domain help the model capture global patterns, which improves the detection accuracy. However, these operations can sometimes lead to misalignment between the semantic and spatial information of different feature maps. To handle this problem, we designed the SAC module, which aligns features from different fusion paths and enhances overall detection performance. The ablation study in Table III demonstrates the effectiveness of these modules.
@@ V Conclusion
This paper proposes UAV-DETR, a real-time end-to-end object detector specifically designed for UAV imagery. By developing the MSFF-FE module, FD module, and SAC module, UAV-DETR helps alleviate the difficulties of detecting small and occluded objects in aerial images. Each of our modules plays a crucial role. The MSFF-FE module enhances multi-scale feature fusion while preserving high-frequency components, the FD module focuses on maintaining spatial details during downsampling, and the SAC module ensures semantic alignment across different feature paths. Experimental results on the VisDrone and UAVVaste datasets demonstrate that our method achieves higher accuracy than existing approaches with similar computational costs while maintaining real-time inference speeds. Future work will focus on improving its robustness to noise.